### v1

# Sparse Approximation in Lattices and Semigroups

Abstract. Given an integer or a non-negative integer solution $\bm{x}$ to a system $\bm{A}\bm{x}=\bm{b}$, where the number of non-zero components of $\bm{x}$ is at most $n$. This paper addresses the following question: How closely can we approximate $\bm{b}$ with $\bm{A}\bm{y}$, where $\bm{y}$ is an integer or non-negative integer solution constrained to have at most $k$ non-zero components with $k<n$? We establish upper and lower bounds for this question in general. In specific cases, these bounds match. The key finding is that the quality of the approximation increases exponentially as $k$ goes to $n$.
1 Introduction
[…]
$\operatorname{app}_{X,k,\|\cdot\|}(\bm{A}):=\max_{\bm{b}\in\bm{A}\cdot X}\min\left\{\|\bm{A}\bm{x}-\bm{b}\|:\bm{x}\in X,\ \left|\operatorname{supp}(\bm{x})\right|\leq k\right\},$
[…]
When $X\in\{\mathbb{Z}^{n},\mathbb{Z}_{\geq 0}^{n}\}$, the task of determining $\operatorname{app}_{X,k,\|\cdot\|}(\bm{A})$ can be viewed as a discrete version of a sparse recovery problem in the context of signal processing. One typical high-level question is to find conditions on $\bm{A}$ such that one can replace the constraint $\left|\operatorname{supp}(\bm{x})\right|\leq k$ by $\|x\|_{1}\leq k$ when $k\in\mathcal{O}(\log_{2}(m))$. The fundamental assumption on $\bm{A}$ is that one knows that a sparse solution exists, or even that it is unique, and one tries to recover it by applying a tractable algorithm. This line of research is very active and there is a huge amount of literature; see for instance [4, 5, 6, 7, 10]. We also refer to the book by Foucart and Rauhut [12] for more information. The direction that we take here deviates from the research questions addressed in the field of sparse recovery in signal processing. In particular, we have no a-priori knowledge about the existence of a sparsest solution.
[…]
Let us now turn to the focus of this paper where $X\in\{\mathbb{Z}^{n},\mathbb{Z}^{n}_{\geq 0}\}$. A first natural question is to ask for an upper bound on $k$ that guarantees $\operatorname{app}_{X,k,\|\cdot\|}(\bm{A})=0$. When $X=\mathbb{Z}^{n}$, such an upper bound on $k$ has been established in [1, Theorem 1]: Let
$\delta(\bm{A}):=\min\{\left|\det\bm{B}\right|:\bm{B}\text{ is an invertible }m\times m\text{ submatrix of }\bm{A}\}.$ (1)
Then the bound in [1, Theorem 1] gives essentially [^1] The result in [1, Theorem 1] is more refined and incorporates the prime factorization of the involved determinants. $\operatorname{app}_{\mathbb{Z}^{n},k,\|\cdot\|_{\infty}}(\bm{A})=0$ whenever $k\geq m+\log_{2}(\delta(\bm{A}))$. We show a result that is more general as it states that $\operatorname{app}_{\mathbb{Z}^{n},k,\|\cdot\|_{\infty}}(\bm{A})\leq\delta(\bm{A})/2^{k-m+1}$, in fact, for $k\geq m+\log_{2}(\delta(\bm{A}))$ we obtain $\operatorname{app}_{\mathbb{Z}^{n},k,\|\cdot\|_{\infty}}(\bm{A})=0$ and thus rediscover the bound in [1, Theorem 1].
In the case $X=\mathbb{Z}^{n}_{\geq 0}$, the analysis of the quantity $\operatorname{app}_{\mathbb{Z}^{n}_{\geq 0},k,\|\cdot\|}(\bm{A})$ is significantly more involved. As before, let us first summarize the results on an upper bound on $k$ that guarantees that $\operatorname{app}_{X,k,\|\cdot\|}(\bm{A})=0$. It has been shown in [11] that $\operatorname{app}_{X,k,\|\cdot\|}(\bm{A})=0$ whenever $k\geq 2m\log_{2}(4m\|\bm{A}\|_{\infty})$. This bound has been improved to $2m\log_{2}(2\sqrt{m}\|\bm{A}\|_{\infty})$ in [2]. Similiar bounds are available when the parameter $\|\bm{A}\|_{\infty}$ is replaced by $\sqrt{\det\bm{A}\bm{A}^{\top}}$; see [2] and [1] for further details.
To the best of our knowledge there are no studies of the approximation parameter $\operatorname{app}_{\mathbb{Z}_{\geq 0}^{n},k,\|\cdot\|}(\bm{A})$ for general values of $k$ available. In the special case when $\operatorname{pos}\{\bm{a}_{1},\ldots,\bm{a}_{n}\}=\mathbb{R}^{m}$ with $\operatorname{pos}Y$ denoting the set of all non-negative combinations of elements in $Y\subseteq\mathbb{R}^{m}$, it turns out that the question of determining $\operatorname{app}_{\mathbb{Z}^{n}_{\geq 0},k,\|\cdot\|}(\bm{A})$ can be tackled similar to the lattice setting where $X=\mathbb{Z}^{n}$. To the contrary, if $\operatorname{pos}\{\bm{a}_{1},\ldots,\bm{a}_{n}\}$ is a pointed cone, we fix a basis $\bm{B}$ consisting of linearly independent column vectors of $\bm{A}$, without loss of generality $\bm{a}_{1},\ldots,\bm{a}_{m}$, and further assume that $\bm{a}_{i}\in\operatorname{pos}\{\bm{a}_{1},\ldots,\bm{a}_{m}\}$ for each $i\in[n]$. For our analysis it will be convenient to work with the norm $\|\cdot\|_{P(\bm{B})}$ induced by the symmetric parallelepiped $P(\bm{B}):=\bm{B}\cdot[-1,1]^{m}$. The dependence of our approximation bound on the matrix $\bm{A}$ is given by the parameter $\mu:=\max_{i\in[n]}\|\bm{a}_{i}\|_{P(\bm{B})}$, which relates the size of $\bm{a}_{i}$ with the size of $\bm{a}_{1},\ldots,\bm{a}_{m}$. We will show that $\operatorname{app}_{\mathbb{Z}^{n}_{\geq 0},k,\|\cdot\|}(\bm{A})$ decreases exponentially in $k$. The precise dependence of $\operatorname{app}_{\mathbb{Z}_{\geq 0}^{n},k,\|\cdot\|}(\bm{A})$ on the parameters $m,k,\mu$, and $\bm{A}$ is presented in Theorem 3.5.
2 Lattices
Let $X=\mathbb{Z}^{n}$ and $\bm{A}\in\mathbb{Z}^{m\times n}$ have full row rank. The set $\bm{A}\cdot X=\bm{A}\mathbb{Z}^{n}$ is the lattice generated by the columns of $\bm{A}$. We refer the reader to [15] for more about lattices, sublattices, and their determinants. Our aim is to find sparse approximations of lattice vectors in $\bm{A}\mathbb{Z}^{n}$. We measure the approximation error with respect to the $\ell_{\infty}$-norm. Observe that if $k\geq m$, we can always express $\bm{b}$ using a basis of the lattice $\bm{A}\mathbb{Z}^{n}$, which consists of $m$ elements. However, it is not necessarily given that such a basis exists among the columns of $\bm{A}$. Even stronger, it could be that no proper subset of the columns is a generating set for the lattice $\bm{A}\mathbb{Z}^{n}$. In this situation, our approximation result below applies. Recall from (1) that $\delta(\bm{A})$ denotes the smallest subdeterminant in absolute value of an invertible $m\times m$ submatrix of $\bm{A}$.
Theorem 2.1
Let $\bm{A}\in\mathbb{Z}^{m\times n}$ have full row rank and $m\leq k\leq n$. Then there exists a $m\times k$ submatrix $\bm{D}$ of $\bm{A}$ such that
$\operatorname{app}_{\mathbb{Z}^{n},k,\|\cdot\|_{\infty}}(\bm{A})\leq\max_{\bm{b}\in\bm{A}\mathbb{Z}^{n}}\min_{\bm{x}\in\mathbb{Z}^{k}}\|\bm{D}\bm{x}-\bm{b}\|_{\infty}\leq\frac{1}{2^{k-m+1}}\cdot\delta(\bm{A}).$
The first inequality holds by definition. To show the second inequality, we rely on the fact that proper subsets of the columns of $\bm{A}$ still generate some sublattice of $\bm{A}\mathbb{Z}^{n}$. Our point of departure is the following lemma that, in particular, holds for $m\times l$ submatrices of $\bm{A}$ with full row rank.
Lemma 1
Let $\bm{D}\in\mathbb{Z}^{m\times l}$ have full row rank and $\Lambda=\bm{D}\mathbb{Z}^{l}$. Then
$\min_{\bm{x}\in\mathbb{Z}^{l}}\|\bm{D}\bm{x}-\bm{d}\|_{\infty}\leq\frac{1}{2}\cdot\det\Lambda$
for all $\bm{d}\in\mathbb{R}^{m}$.
Proof
There exists a unimodular matrix $\bm{U}\in\mathbb{Z}^{l\times l}$ such that $\bm{D}\bm{U}=[\bm{B},\bm{0}]$ is in Hermite normal form; cf. [18, Chapter 4]. In particular, the matrix $\bm{B}$ is a lower triangular matrix and $\det\Lambda=\bm{B}_{1,1}\cdot\ldots\cdot\bm{B}_{n,n}$. Define $\bm{z}\in\mathbb{Z}^{m}$ recursively such that
$\bm{z}_{i}:=\left\lceil\frac{\bm{d}_{i}-\sum_{j=1}^{i-1}\bm{B}_{i,j}\bm{z}_{j}}{\bm{B}_{i,i}}\right\rfloor,$
where $\left\lceil x\right\rfloor$ denotes the integer closest to $x\in\mathbb{R}$. Then we have
$\left|(\bm{D}\bm{U}\bm{z}-\bm{d})_{i}\right|=\left|(\bm{B}\bm{z}-\bm{d})_{i}\right|=\left|\bm{B}_{i,i}\left(\frac{\sum_{j=1}^{i-1}\bm{B}_{i,j}\bm{z}_{j}-\bm{d}_{i}}{\bm{B}_{i,i}}+\bm{z}_{i}\right)\right|\leq\tfrac{1}{2}\bm{B}_{i,i}$
for all $i\in[m]$. The last term above is upper bounded by $1/2\cdot\det\Lambda$. So the integer vector $\bm{U}\bm{z}$ satisfies the claimed bound. ∎
Proof (of Theorem 2.1)
As mentioned, the first inequality holds by definition. We prove the second inequality by recursively constructing submatrices $\bm{D}$ of $\bm{A}$ to which Lemma 1 applies. Let $\bm{B}$ be a $m\times m$ submatrix of $\bm{A}$ such that $\left|\det\bm{B}\right|=\delta(\bm{A})$. We set $\bm{D}_{0}:=\bm{B}$ and $\Lambda_{0}:=\bm{D}_{0}\mathbb{Z}^{m}$. For $i\in[n-m]$, choose a column of $\bm{A}$, say $\bm{v}_{i}$, such that $\bm{v}_{i}\notin\Lambda_{i-1}$. If there is no such column, we simply select an arbitrary column. Then we define $\bm{D}_{i}:=[\bm{D}_{i-1},\bm{v}_{i}]\in\mathbb{Z}^{m+i}$ and $\Lambda_{i}:=\bm{D}_{i}\mathbb{Z}^{m+i}$. By Lemma 1, we have
$\min_{\bm{x}\in\mathbb{Z}^{m+i}}\|\bm{D}_{i}\bm{x}-\bm{b}\|_{\infty}\leq\frac{1}{2}\cdot\det\Lambda_{i}.$ (2)
for all $i\in\{0,\ldots,n-m\}$ and $\bm{b}\in\bm{A}\mathbb{Z}^{n}$. To finish the proof, we analyze $\det\Lambda_{i}$ for $i\in\{0,\ldots,n-m\}$. If $i=0$, we get $\det\Lambda_{0}=\left|\det\bm{B}\right|=\delta(\bm{A})$ and the claim follows. Fix some $i\in[n-m]$ and recall that $\bm{v}_{i}$ is the column that we added in the $i$-th step. If $\bm{v}_{i}\in\Lambda_{i-1}$, our construction implies that all columns of $\bm{A}$ that are not columns of $\bm{D}_{i-1}$ are contained in $\Lambda_{i-1}$. It follows that $\det\bm{A}\mathbb{Z}^{n}=\det\Lambda_{i-1}=\det\Lambda_{i}$ and $\min_{\bm{x}\in\mathbb{Z}^{m+i}}\|\bm{D}_{i}\bm{x}-\bm{b}\|_{\infty}=0$ for all $\bm{b}\in\bm{A}\mathbb{Z}^{n}$, which implies the claim. So suppose that $\bm{v}_{i}\notin\Lambda_{i-1}$. This is equivalent to $\Lambda_{i-1}\subsetneq\Lambda_{i}$. Since $\Lambda_{i-1}$ is a proper sublattice of $\Lambda_{i}$, we obtain that $\det\Lambda_{i}\neq\det\Lambda_{i-1}$ and $\det\Lambda_{i}$ divides $\det\Lambda_{i-1}$. Thus, we have $\det\Lambda_{i}\leq 1/2\cdot\det\Lambda_{i-1}$. By recursively applying this inequality, we recover
$\det\Lambda_{i}\leq\frac{1}{2^{i}}\cdot\det\Lambda_{0}=\frac{1}{2^{i}}\cdot\delta(\bm{A}).$
Plugging this into (2) and choosing $i=k-m$ proves the claim. ∎
From Theorem 2.1 we conclude that the approximation factor drops exponentially in $k$. One can refine our argument from above by choosing the $\bm{v}_{i}$ that minimizes $\det\Lambda_{i}$ when constructing $\bm{D}$. This relates to incorporating the prime factorization of $\delta(\bm{A})$ similar to the techniques in the proof of [1, Theorem 1]. We omit this here for the sake of brevity and since we are mainly interested in proving the exponential behavior in $k$. The bound in Theorem 2.1 also implies bounds on the exact representation of vectors in lattices in terms of some generating set $\bm{A}$: For $k\geq m+\log_{2}(\delta(\bm{A}))$ we obtain $\operatorname{app}_{\mathbb{Z}^{n},k,\|\cdot\|_{\infty}}(\bm{A})\leq 1/2$. Since $\operatorname{app}_{\mathbb{Z}^{n},k,\|\cdot\|_{\infty}}(\bm{A})$ is an integer, we conclude that $\operatorname{app}_{\mathbb{Z}^{n},k,\|\cdot\|_{\infty}}(\bm{A})=0$. This is essentially the order of magnitude presented in [1, Theorem 1], which is not a coincidence as our techniques extend the approach used in [1].
3 Semigroups
Let $X=\mathbb{Z}^{n}_{\geq 0}$ and $\bm{A}=(\bm{a}_{1},\ldots,\bm{a}_{m})\in\mathbb{Z}^{m\times n}$ have full row rank. Then $\bm{A}\cdot X$ is the semigroup generated by the columns of $\bm{A}$. If $\operatorname{pos}\{\bm{a}_{1}\ldots,\bm{a}_{n}\}=\mathbb{R}^{m}$, each subset $I\subseteq[n]$ with $\operatorname{pos}\{\bm{a}_{i}:i\in I\}=\mathbb{R}^{m}$ generates non-negatively the lattice spanned by $\bm{a}_{i}$ for $i\in I$. Therefore, one can apply the lattice technique presented in the previous section. So we suppose that $\operatorname{pos}\{\bm{a}_{1},\ldots,\bm{a}_{n}\}$ generates a pointed cone. We begin our analysis with the case when $m=1$. The assumption that $\bm{a}_{1},\ldots,\bm{a}_{n}$ generate a pointed cone implies without loss of generality that $0<\bm{a}_{1}\leq\bm{a}_{2}\leq\ldots\leq\bm{a}_{n}$. Our goal is to determine the approximation error with respect to the absolute value, that is, $\operatorname{app}_{\bm{a}^{\top}\mathbb{Z}^{n},k,|\cdot|}(\bm{a})$ for a fixed $k\geq 1$. We abbreviate this parameter by $\operatorname{app}_{k}(\bm{a})$. To bound $\operatorname{app}_{k}(\bm{a})$, we apply the following strategy: We fix the smallest integer $\bm{a}_{1}$ and bound the slightly refined parameter
$\operatorname{app}_{\bm{a}_{1},k}(\bm{a}):=\max_{b\in\bm{a}^{\top}\mathbb{Z}^{n}_{\geq 0}}\min\left\{\left|\bm{a}^{\top}\bm{x}-b\right|:\bm{x}\in\mathbb{Z}^{n}_{\geq 0},\ \left|\operatorname{supp}(\bm{x})\backslash\{1\}\right|\leq k-1\right\},$
which corresponds to the best approximation that uses $\bm{a}_{1}$. By definition, we have $\operatorname{app}_{k}(\bm{a})\leq\operatorname{app}_{\bm{a}_{1},k}(\bm{a})$. Using the pigeonhole principle one can derive an upper bound on $\operatorname{app}_{n-1}(\bm{a})$. Note that this bound applies to real valued $\bm{a}$.
Theorem 3.1
Let $\bm{a}\in\mathbb{R}^{n}$ with $0<\bm{a}_{1}\leq\ldots\leq\bm{a}_{n}$. Then we have
$\operatorname{app}_{n-1}(\bm{a})\leq\operatorname{app}_{\bm{a}_{1},n-1}(\bm{a})\leq\frac{1}{2^{n-1}}\cdot\bm{a}_{1}.$
Next we apply Theorem 3.1 repeatedly to obtain a bound for general values of $k$. This bound shows that the approximation factor decreases exponentially as $k$ goes to $n$.
Theorem 3.2
Let $\bm{a}\in\mathbb{R}^{n}$ with $0<\bm{a}_{1}\leq\ldots\leq\bm{a}_{n}$ and $1\leq k\leq n$. Then we have
$\operatorname{app}_{k}(\bm{a})\leq\operatorname{app}_{\bm{a}_{1},k}(\bm{a})\leq\left(\frac{1}{2^{k-1}}-\frac{1}{2^{n-1}}\right)\cdot\bm{a}_{1}.$
Theorems 3.1 and 3.2 under the assumption that $\bm{a}\in\mathbb{Z}^{n}$ follow from their extensions to general values of $m$ that we present below.
Let us first comment on the question whether the bounds in Theorems 3.1 and 3.2 are tight. The bound in Theorem 3.1 is tight, even when $\bm{a}$ is integral, see Proposition 1 in Section 6. For $k=1$, Theorem 3.2 can be strengthened to $\operatorname{app}_{\bm{a}_{1},1}(\bm{a})\leq 1/2\cdot\bm{a}_{1}$, based on the observation that we can write every integer as an integer multiple of $\bm{a}_{1}$ plus or minus a remainder of size at most $1/2\cdot\bm{a}_{1}$. This construction is tight. These observations for $k\in\{1,n-1\}$ suggest that $\operatorname{app}_{k}(\bm{a})\leq 1/2^{k}\cdot\bm{a}_{1}$ might be the correct bound for all $1\leq k\leq n$. This is, however not true, even when $k=2$. In fact Theorem 3.3 below presents a tight bound for $k=2$; see also Proposition 2 in Section 6. To prepare for the presentation of this bound, let us define the sequence $t_{0}:=1$ and $t_{i}:=t_{i-1}\cdot(t_{i-1}+1)$ for $i\geq 1$. This sequence $\{t_{i}\}_{i\in\mathbb{N}}$ is well-known and appears in several papers related to combinatorial optimization problems; see, for instance, [13, 14, 19]. Moreover, the sequence $\{t_{i}+1\}_{i\in\mathbb{N}}$ is known as Sylvester’s sequence. It gives among others the optimal solution to approximating $1$ by egyptian fractions from below; see [8, 16, 17, 20] for some appearances of Sylvester’s sequence in the literature. For $m\in\mathbb{N}$, we define
$\varphi(m):=\sum_{i=1}^{m}\frac{1}{t_{i}}.$
Theorem 3.3
Let $\bm{a}\in\mathbb{R}^{n}$ with $0<\bm{a}_{1}\leq\ldots\leq\bm{a}_{n}$. Then we have
$\operatorname{app}_{2}(\bm{a})\leq\operatorname{app}_{\bm{a}_{1},2}(\bm{a})\leq\frac{\varphi(n-2)}{2\varphi(n-2)+1}\cdot\bm{a}_{1}.$
Note that for $n=3$ we get an upper bound of $1/4\cdot\bm{a}_{1}$, for $n=4$ we get $2/7\cdot\bm{a}_{1}$ and $29/100\cdot\bm{a}_{1}$ for $n=5$. In the limit, we obtain $\lim_{m\to\infty}\varphi(m)\approx 0.691$ which gives us a bound of approximately $0.2901\cdot\bm{a}_{1}$. So we are slightly worse than the bound $1/2^{k}\cdot\bm{a}_{1}=1/4\cdot\bm{a}_{1}$ for $k=2$. We prove Theorem 3.3 in Section 5.
Let us now turn to the extensions of Theorem 3.1 and 3.2 to general values of $m\geq 2$. Our strategy to extend these results is motivated by the analysis when $m=1$: We fix an invertible $m\times m$ submatrix $\bm{B}$ of $\bm{A}$. Without loss of generality let $\bm{B}=(\bm{a}_{1},\ldots,\bm{a}_{m})$. In addition, we assume that $\bm{a}_{i}\in\operatorname{pos}\{\bm{a}_{1},\ldots,\bm{a}_{m}\}$ for each $i\in[n]$. This assumption resembles the case $m=1$, where we fixed $\bm{B}=\bm{a}_{1}$ and also assumed that $\bm{a}_{i}\in\operatorname{pos}\{\bm{a}_{1}\}$. We measure our approximation error in terms of the symmetric parallelepiped spanned by $\bm{a}_{1},\ldots,\bm{a}_{m}$, which is defined by $P(\bm{B}):=\bm{B}\cdot[-1,1]^{m}$. Our errors will be taken with respect to the norm $\|\cdot\|_{P(\bm{B})}$, the norm with unit ball $P(\bm{B})$. This also is in correspondence with what we did when $m=1$ because, if $m=1$, we have $\bm{B}=\bm{a}_{1}$ and the rescaled norm induced by $1/\bm{a}_{1}\cdot P(\bm{a}_{1})=[-1,1]$ equals simply the absolute value. As before, let us abbreviate $\operatorname{app}_{\bm{A}\mathbb{Z}^{n}_{\geq 0},k,\|\cdot\|_{P(\bm{B})}}(\bm{A})$ by $\operatorname{app}_{k}(\bm{A})$ for a fixed value of $k\geq 1$. We are now prepared to investigate the refined parameter
$\operatorname{app}_{\bm{B},k}(\bm{A}):=\max_{\bm{b}\in\bm{A}\mathbb{Z}^{n}_{\geq 0}}\min\left\{\|\bm{A}\bm{x}-\bm{b}\|_{P(\bm{B})}:\bm{x}\in\mathbb{Z}^{n}_{\geq 0},\ \left|\operatorname{supp}(\bm{x})\backslash[m]\right|\leq k-m\right\}.$
Theorem 3.4
Let $\bm{A}=(\bm{a}_{1},\ldots,\bm{a}_{n})\in\mathbb{Z}^{m\times n}$ have full row rank, $m\leq n-1$, $\bm{B}=(\bm{a}_{1},\ldots,\bm{a}_{m})$ be invertible, $\mu=\max_{i\in[n]}\|\bm{a}_{i}\|_{P(\bm{B})}$, and $\bm{a}_{i}\in\operatorname{pos}\{\bm{a}_{1},\ldots,\bm{a}_{m}\}$ for all $i\in[n]$. Then we have
$\operatorname{app}_{n-1}(\bm{A})\leq\operatorname{app}_{\bm{B},n-1}(\bm{A})\leq\frac{1}{2^{\frac{n-m}{m}}}\cdot\mu^{\frac{m-1}{m}}\cdot\left|\det\bm{B}\right|^{\frac{m-1}{m}}.$
Theorem 3.4 implies Theorem 3.1 when $m=1$ and $\bm{a}\in\mathbb{Z}^{n}$. As for $m=1$, Theorem 3.4 applied repeatedly allows us to observe that the approximation factor decreases exponentially as $k$ goes to $n$.
Theorem 3.5
Let $\bm{A}=(\bm{a}_{1},\ldots,\bm{a}_{n})\in\mathbb{Z}^{m\times n}$ have full row rank. Let $\bm{B}=(\bm{a}_{1},\ldots,\bm{a}_{m})$ be invertible, $\mu=\max_{i\in[n]}\|\bm{a}_{i}\|_{P(\bm{B})}$, and $\bm{a}_{i}\in\operatorname{pos}\{\bm{a}_{1},\ldots,\bm{a}_{m}\}$ for all $i\in[n]$. Then we have
$\operatorname{app}_{k}(\bm{A})\leq\operatorname{app}_{\bm{B},k}(\bm{A})\leq\frac{1}{2^{\frac{1}{m}}-1}\cdot\left(\frac{1}{2^{\frac{k-m}{m}}}-\frac{1}{2^{\frac{n-m}{m}}}\right)\cdot\mu^{\frac{m-1}{m}}\cdot\left|\det\bm{B}\right|^{\frac{m-1}{m}}$
for all $m\leq k\leq n$.
By evaluating the inequality in Theorem 3.5 at $m=1$ and rescaling the norm, we recover the bound given in Theorem 3.2 provided that $\bm{a}\in\mathbb{Z}^{n}$.
A few comments on the bound in Theorem 3.5 are in order. Notice that $\operatorname{app}_{k}(\bm{A})\leq\operatorname{app}_{\bm{B},k}(\bm{A})\leq 1/2$ is a trivial upper bound that holds for all $k\geq m$. For small values of $k$, the bound in Theorem 3.5 is worse than $1/2$. In particular, one can calculate that the bound in Theorem 3.5 improves upon $1/2$ if we have $k\geq 2m+m\log_{2}(cm\mu^{(m-1)/m}\left|\det\bm{B}\right|^{(m-1)/m})$ for some constant $c>0$. These calculations utilize the fact that $1/(2^{1/m}-1)=\mathcal{O}(m)$ for $m\geq 1$. However, recall that a sparsity result in [11, Theorem 1 (ii)] implies that we have an exact sparse representation using only $2m\log_{2}(4m\|\bm{A}\|_{\infty})$ vectors. Adapting the argument in [11] to the parameters $\mu$ and $\left|\det\bm{B}\right|$ one can derive that there is always an exact sparse representation using $2m\log_{2}(4m\mu\left|\det\bm{B}\right|^{1/m})$ vectors. Hence, our bounds are relevant whenever we have
$2m+m\log_{2}(cm\mu^{\frac{m-1}{m}}\left|\det\bm{B}\right|^{\frac{m-1}{m}})\leq k\leq 2m\log_{2}\left(4m\mu\left|\det\bm{B}\right|^{\frac{1}{m}}\right).$
This is the case when $\mu$ is relatively large compared to $m$ and $\left|\det\bm{B}\right|$. Similar to the case $X=\mathbb{Z}^{n}$, the results presented above imply bounds on the size of an exact representation when $X=\mathbb{Z}^{n}_{\geq 0}$. This follows from Theorem 3.4.
Corollary 1
Let $\bm{A}=(\bm{a}_{1},\ldots,\bm{a}_{n})\in\mathbb{Z}^{m\times n}$ have full row rank. Let $\bm{B}=(\bm{a}_{1},\ldots,\bm{a}_{m})$ be invertible, $\mu=\max_{i\in[n]}\|\bm{a}_{i}\|_{P(\bm{B})}$, and $\bm{a}_{i}\in\operatorname{pos}\{\bm{a}_{1},\ldots,\bm{a}_{n}\}$ for all $i\in[n]$. Then we have $\operatorname{app}_{n-1}(\bm{A})=0$ for
$n\geq m+m\log_{2}(\mu^{(m-1)/m}\left|\det\bm{B}\right|^{(2m-1)/m}).$
When $m=1$, we recover a tight bound on the sparse representation of integers in semigroups; cf. [1, Theorem 3]. For $m\geq 2$, the bound in Corollary 1 outperforms the bound $2m\log_{2}(4m\mu\left|\det\bm{B}\right|^{1/m})$ in [11] whenever $\mu$ is sufficiently large.
It is important to point out that our bounds for $m\geq 2$ depend on the factor $\mu$ whereas the bounds for $m=1$ are independent of $\mu$. This is not a coincidence. In fact, for $m\geq 2$ Proposition 3 in Section 6 demonstrates that a dependence on $\mu$ (or some related parameter measuring the size of $\bm{A}$) is necessary to obtain approximation factors that decrease in $k$.
Finally, let us mention that our upper bounds are based on the assumption that $\bm{a}_{i}\in\operatorname{pos}\{\bm{a}_{1},\ldots,\bm{a}_{m}\}$ for all $i\in[n]$. A closer examination of the techniques in Section 4 reveals that this assumption is not necessary. However, eliminating it requires us to introduce an additional factor of $2^{(m-1)/m}$ in Theorem 3.4 and Theorem 3.5. We state this here without a proof.
Remark 1
Let $\bm{A}=(\bm{a}_{1},\ldots,\bm{a}_{n})\in\mathbb{Z}^{m\times n}$ with $m\leq k\leq n$ have full row rank, $\bm{B}=(\bm{a}_{1},\ldots,\bm{a}_{m})$ be invertible and $\mu=\max_{i\in[n]}\|\bm{a}_{i}\|_{P(\bm{B})}$. Then we have
$\operatorname{app}_{k}(\bm{A})\leq\operatorname{app}_{\bm{B},k}(\bm{A})\leq\frac{2^{\frac{m-1}{m}}}{2^{\frac{1}{m}}-1}\cdot\left(\frac{1}{2^{\frac{k-m}{m}}}-\frac{1}{2^{\frac{n-m}{m}}}\right)\cdot\mu^{\frac{m-1}{m}}\cdot\left|\det\bm{B}\right|^{\frac{m-1}{m}}.$
The proof techniques will be presented in the next section. In fact, they extend to arbitrary real matrices. This requires further technicalities that we refrain from discussing in the paper.
4 Proofs of Theorem 3.4, Theorem 3.5, and Corollary 1
Let us start to describe the major steps required to prove Theorem 3.4. The first part follows a similar approach to the techniques in [11]: We fix some $\bm{b}\in\bm{A}\cdot\mathbb{Z}^{n}_{\geq 0}$. This vector has a representation of the form $\bm{b}=\lambda_{1}\bm{a}_{1}+\ldots+\lambda_{n}\bm{a}_{n}$ with $\lambda_{i}\in\mathbb{N}$ for all $i\in[n]$. We tile $\operatorname{pos}\{\bm{a}_{1},\ldots,\bm{a}_{m}\}$ with sufficiently small copies of the parallelepiped spanned by $\bm{a}_{1},\ldots,\bm{a}_{m}$. Then we count how many subsums of $\bm{b}=\lambda_{1}\bm{a}_{1}+\ldots+\lambda_{n}\bm{a}_{n}$ are contained in the copies of the sufficiently small parallelepipeds. It turns out that the number of subsums grows exponentially in $n$ and the number of parallelepipeds that we need to consider only grows exponentially in $m$. Once this has been established one could derive a first bound that is, however, weaker than the one stated in Theorem 3.4. (The key distinction is that one would replace the expression $\mu^{(m-1)/m}\cdot\left|\det\bm{B}\right|^{(m-1)/m}$ by the expression $\mu\cdot\left|\det\bm{B}\right|$.) To get Theorem 3.4, a second ingredient must be incorporated, one that, to the best of our knowledge, has not previously been utilized: the idea is to show that we only need to consider so-called incomparable subsums.
To make this notion precise, let $(P_{m,s},\preceq)$ be the partially order set given by $P_{m,s}:=\mathbb{Z}^{m}\cap[0,s]^{m}$ and $\bm{y}\preceq\bm{z}$ if and only if $\bm{y}_{1}\leq\bm{z}_{1},\ldots,\bm{y}_{m}\leq\bm{z}_{m}$ for $m,s\geq 1$. An antichain of $(P_{m,s},\preceq)$ is a subset of $P_{m,s}$ whose elements are pairwise incomparable. An upper bound on the size of a largest antichain in $(P_{m,s},\preceq)$ can be established easily.
Lemma 2
Let $m,s\in\mathbb{N}_{\geq 1}$. The largest antichain of $(P_{m,s},\preceq)$ is bounded above by $(s+1)^{m-1}$. It is in the order of $\Theta(s^{m-1})$.
Proof
We determine the upper bound first. By Dilworth’s theorem [9], the largest antichain equals the minimal size of a chain cover. Thus, we can construct a chain cover to get an upper bound on the size of the largest antichain. Define the chains $C_{\bm{z}_{1},\ldots,\bm{z}_{m-1}}:=\{(\bm{z}_{1},\ldots,\bm{z}_{m-1},0)^{\top},\ldots,(\bm{z}_{1},\ldots,\bm{z}_{m-1},s)^{\top}\}$ for fixed values $\bm{z}_{1},\ldots,\bm{z}_{m-1}\in\{0,\ldots,s\}$. Those chains cover $P_{m,s}$. In total, there are $(s+1)^{m-1}$ chains of the type $C_{\bm{z}_{1},\ldots,\bm{z}_{m-1}}$. This gives the desired upper bound.
For the asymptotic bound, note that the upper bound implies that the size of a largest antichain is in $\mathcal{O}(s^{m-1})$. Hence, it suffices to construct an antichain of size $\Omega(s^{m-1})$. Define the subset of $P_{m,s}$ given by $S:=\{\bm{z}\in P_{m,s}:\sum_{i=1}^{m}\bm{z}_{i}=s\}$. The elements in $S$ are incomparable by construction. From Ehrhart theory, cf. [3, Chapter 3], it follows that $\left|S\right|=\Omega(s^{m-1})$ since the elements in $S$ lie on an $(n-1)$-dimensional affine hyperplane. ∎
Let $\operatorname{par}\{\bm{a}_{1},\ldots,\bm{a}_{m}\}$ denote the half-open parallelepiped spanned by the basis vectors $\bm{a}_{1},\ldots,\bm{a}_{m}$. Taking the parameter $\mu$ into consideration one has that
$\bm{a}_{i}\in\mu\cdot\operatorname{par}\{\bm{a}_{1},\ldots,\bm{a}_{m}\}\text{ for all }i\in[n].$
Proof (of Theorem 3.4)
Let $\bm{b}:=\lambda_{1}\bm{a}_{1}+\ldots+\lambda_{n}\bm{a}_{n}$ with $\lambda_{i}\in\mathbb{N}$ for all $i\in[n]$. Since we fix $\bm{a}_{1},\ldots,\bm{a}_{m}$ in our non-negative integral combination, we assume without loss of generality $\lambda_{i}=0$ for $i\in[m]$. So we have $\bm{b}=\lambda_{m+1}\bm{a}_{m+1}+\ldots+\lambda_{n}\bm{a}_{n}$. Let us abbreviate $\lambda:=\lambda_{m+1}+\ldots+\lambda_{n}$. We have that $\bm{b}\in\mu\cdot\lambda\cdot\operatorname{par}\{\bm{a}_{1},\ldots,\bm{a}_{m}\}$ as $\bm{a}_{i}\in\mu\cdot\operatorname{par}\{\bm{a}_{1},\ldots,\bm{a}_{m}\}$ for all $i\in[n]$. We tile $\mu\cdot\lambda\cdot\operatorname{par}\{\bm{a}_{1},\ldots,\bm{a}_{m}\}$ using $\mu^{m}\cdot\lambda^{m}$ copies of the parallelepiped $\operatorname{par}\{\bm{a}_{1},\ldots,\bm{a}_{m}\}$. We refine this tiling by further dividing each $\operatorname{par}\{\bm{a}_{1},\ldots,\bm{a}_{m}\}$ into $l^{m}$ parellelepipeds of size and shape $1/l\cdot\operatorname{par}\{\bm{a}_{1},\ldots,\bm{a}_{m}\}$. In total, we have $\mu^{m}\cdot\lambda^{m}\cdot l^{m}$ parallelepipeds tiling $\mu\cdot\lambda\cdot\operatorname{par}\{\bm{a}_{1},\ldots,\bm{a}_{m}\}$. Let $\tilde{\bm{a}}_{i}:=\lambda_{i}\bm{a}_{i}$. Suppose that two subsums of $\tilde{\bm{a}}_{m+1}+\ldots+\tilde{\bm{a}}_{n}$ are in the same tile, up to adding multiples of $\bm{a}_{1},\ldots,\bm{a}_{m}$ to one of the subsums. In other words, suppose that we have two different sets $I,J\subseteq\{m+1,\ldots,n\}$ such that $\mu_{1}\bm{a}_{1}+\ldots+\mu_{m}\bm{a}_{m}+\sum_{i\in I}\tilde{\bm{a}}_{i}$ and $\sum_{j\in J}\tilde{\bm{a}}_{j}$ are in the same tile for $\mu_{1},\ldots,\mu_{m}\in\mathbb{N}$. This implies $\|\mu_{1}\bm{a}_{1}+\ldots+\mu_{m}\bm{a}_{m}+\sum_{i\in I}\tilde{\bm{a}}_{i}-\sum_{j\in J}\tilde{\bm{a}}_{j}\|_{P(\bm{B})}\leq 1/l$. Since $I,J$ are different sets, we assume without loss of generality that there exists $j^{*}\in J$ with $j^{*}\notin I$. If this is not the case, we have $J\subsetneq I$ which implies $\sum_{j\in J}\tilde{\bm{a}}_{j}<\sum_{i\in I}\tilde{\bm{a}}_{i}$. So we have $\mu_{i}=0$ for all $i\in[m]$ and we can interchange the role of $I$ and $J$. Using this, we construct a new approximate vector
$\bar{\bm{b}}:=\bm{b}+\mu_{1}\bm{a}_{1}+\ldots+\mu_{m}\bm{a}_{m}+\sum_{i\in I}\tilde{\bm{a}}_{i}-\sum_{j\in J}\tilde{\bm{a}}_{j}.$
This is an element of the semigroup that is generated by $n-1$ elements since we do not use the generator $\tilde{\bm{a}}_{j^{*}}$. We get
$\|\bar{\bm{b}}-\bm{b}\|_{P(\bm{B})}=\|\mu_{1}\bm{a}_{1}+\ldots+\mu_{m}\bm{a}_{m}+\sum_{i\in I}\tilde{\bm{a}}_{i}-\sum_{j\in J}\tilde{\bm{a}}_{j}\|_{P(\bm{B})}\leq\frac{1}{l}.$ (3)
It remains to find a bound on $l$. To accomplish this task, we investigate how fine we want to tile $\operatorname{par}\{\bm{a}_{1},\ldots,\bm{a}_{m}\}$ while simultaneously ensuring that two subsums of $\tilde{\bm{a}}_{m+1}+\ldots+\tilde{\bm{a}}_{n}$ are in the same tile, up to adding multiples of $\bm{a}_{1},\ldots,\bm{a}_{m}$ to one of the subsums. We have $l^{m}$ tiles in $\operatorname{par}\{\bm{a}_{1},\ldots,\bm{a}_{m}\}$. Fix one of those tiles and call it $T$. We define $T(\mu_{1},\ldots,\mu_{m}):=T+\sum_{i=1}^{n}\mu_{i}\bm{a}_{i}$ for $\mu_{1},\ldots,\mu_{m}\in\{0,\ldots,\mu\cdot\lambda-1\}$. These are the translates of $T$ modulo $\bm{a}_{1},\ldots,\bm{a}_{m}$ that are contained in $\mu\cdot\lambda\cdot\operatorname{par}\{\bm{a}_{1},\ldots,\bm{a}_{m}\}$. There are $\mu^{m}\cdot\lambda^{m}$ of those translates. We claim that if there are more than $\mu^{m-1}\cdot\lambda^{m-1}$ subsums in different $T(\mu_{1},\ldots,\mu_{m})$, we can add multiples of $\bm{a}_{1},\ldots,\bm{a}_{m}$ to one of the subsums and end up in the same translated tile. To see this, we define a partial order on the set $\mathcal{T}$ of the translates of $T(\mu_{1},\ldots,\mu_{m})$ for $\mu_{1},\ldots,\mu_{m}\in\{0,\ldots,\mu\cdot\lambda-1\}$. Let $T(\mu_{1},\ldots,\mu_{m})\preceq T(\mu^{\prime}_{1},\ldots,\mu^{\prime}_{m})$ if and only if $\mu_{1}\leq\mu^{\prime}_{1},\ldots,\mu_{m}\leq\mu^{\prime}_{m}$. This partial order implies that if two subsums are contained in translated tiles that obey the ordering above, we are done since $T(\mu_{1},\ldots,\mu_{m})+\sum_{i=1}^{m}(\mu^{\prime}_{i}-\mu_{i})=T(\mu^{\prime}_{1},\ldots,\mu^{\prime}_{m})$ and $\mu^{\prime}_{i}-\mu_{i}\geq 0$ for all $i\in[m]$. Hence, we want to find the largest subset of $\mathcal{T}$ whose elements are pairwise incomparable with respect to the partial order. This is equivalent to determining the largest antichain in $(\mathcal{T},\preceq)$. Observe that $(\mathcal{T},\preceq)$ is isomorphic to $(P_{m,\mu\cdot\lambda-1},\preceq)$. So we can apply Lemma 2 to obtain the upper bound $\mu^{m-1}\cdot\lambda^{m-1}$ on the size of the largest antichain in $(\mathcal{T},\preceq)$. This holds for all $l^{m}$ tiles. Hence, by the pigeonhole principle we must ensure that
$2^{n-m}\geq\mu^{m-1}\cdot\lambda^{m-1}\cdot l^{m}$
to guarantee that two subsums are, up to adding multiples of $\bm{a}_{1},\ldots,\bm{a}_{m}$, in the same tile. To get the best possible such bound, we choose $l$ to be as large as possible such that the inequality holds. Since we have ignored for simplicity a couple of improving steps in the analysis above (such as the empty subsum ruling out more tiles), we do not round $l$ to the nearest integer, but simply select $l:=2^{n-m}\cdot(\mu\cdot\lambda)^{-(m-1)}$. After plugging this into (3) it remains to show that $\lambda\leq\left|\det\bm{B}\right|$. To verify the latter inequality, we use again the pigeonhole principle.
For the purpose of deriving a contradiction assume that $\lambda\geq\left|\det\bm{B}\right|+1$. We analyze the sum
$\underbrace{\bm{a}_{m+1}+\ldots+\bm{a}_{m+1}}_{\lambda_{m+1}\text{ times }}+\ldots+\underbrace{\bm{a}_{n}+\ldots+\bm{a}_{n}}_{\lambda_{n}\text{times}}.$
Let $\bm{y}_{i}$ denote the subsum of the first $i$ terms from above for $i\in[\sum_{i=m+1}^{n}\lambda_{i}]$. There are at least $\left|\det\bm{B}\right|+1$ such subsums. Moreover, there exist $\left|\det\bm{B}\right|$ different elements of the form $\bm{y}+\bm{B}\mathbb{Z}^{m}$ for $\bm{y}\in\mathbb{Z}^{m}$. By the pigeonhole principle, we can find two different subsums $\bm{y}_{i}$ and $\bm{y}_{j}$ with $i<j$ such that $\bm{y}_{j}-\bm{y}_{i}\in\bm{B}\mathbb{Z}^{m}$ and $\bm{y}_{j}-\bm{y}_{i}\in\operatorname{pos}\{\bm{a}_{1},\ldots,\bm{a}_{m}\}$ as $i<j$. So we have $\bm{y}_{j}-\bm{y}_{i}=\sum_{i=1}^{m}\mu_{i}\bm{a}_{i}$ for non-negative integral values $\mu_{1},\ldots,\mu_{m}$. Instead of approximating $\bm{b}$, we approximate the new integer vector $\bm{b}-(\bm{y}_{j}-\bm{y}_{i})\in\operatorname{pos}\{\bm{a}_{1},\ldots,\bm{a}_{m}\}$. Once we have an approximation of this new vector, we add the vector $\sum_{i=1}^{m}\mu_{i}\bm{a}_{i}$ back and obtain an approximation of $\bm{b}$ with the same error. Since $\bm{b}-(\bm{y}_{j}-\bm{y}_{i})$ has a strictly smaller coordinate sum than $\bm{b}$, one can repeat these steps until the coordinate sum is bounded by $\left|\det\bm{B}\right|$. ∎
The lower bound in Lemma 2 implies that we cannot asymptotically improve upon the bound $\mu^{m-1}\cdot\lambda^{m-1}$ when estimating the size of a largest antichain.
Proof (of Theorem 3.5)
The claim follows immediately when $k=n$. So suppose that $k\leq n-1$. Let ${\bm{b}}=\bm{A}{\bm{x}}$ for ${\bm{x}}\in\mathbb{Z}^{n}_{\geq 0}$ be given. We set $\bm{x}^{0}:=\bm{x}$ and $I_{0}:=\operatorname{supp}(\bm{x})\backslash[m]$. For $j\geq 1$, we define $I_{j}$ and $\bm{x}^{j}$ recursively as follows:
  1. 1.
    Let $\bm{A}_{I_{j-1}}$ be the matrix with columns in $[m]\cup I_{j-1}$ and $\bm{b}^{j-1}=\bm{A}{\bm{x}}^{j-1}$.
  2. 2.
    Apply Theorem 3.4 to $\bm{b}^{j-1}$ and $\bm{A}_{I_{j-1}}$. We obtain an approximation ${\bm{b}}^{j}=\bm{A}\bm{x}^{j}$ for some $\bm{x}^{j}\in\mathbb{Z}^{n}_{\geq 0}$ with $\operatorname{supp}(\bm{x}^{j})\subseteq[m]\cup I_{j-1}$ and $\left|\operatorname{supp}(\bm{x}^{j})\backslash[m]\right|<|I_{j-1}|$.
  3. 3.
    If $\left|\operatorname{supp}(\bm{x}^{j})\backslash[m]\right|\leq k-m$, we are done. Otherwise, define $I_{j}:=\operatorname{supp}(\bm{x}^{j})\backslash[m]$ and return to the first step.
Since we reduce $|I_{j}|$ by at least one in each step, this procedure terminates with some $\bm{x}^{l}$ such that $\left|\operatorname{supp}(\bm{x}^{l})\backslash[m]\right|\leq k-m$ and $1\leq l\leq n-k$. We observe
$\operatorname{app}_{k}(\bm{A})\leq\operatorname{app}_{\bm{B},k}(\bm{A})\leq\sum_{j=0}^{l-1}\operatorname{app}_{\bm{B},\left|I_{j}\right|+m-1}(\bm{A}_{I_{j}}).$ (4)
It remains to bound $\operatorname{app}_{\bm{B},\left|I_{j}\right|+m-1}(\bm{A}_{I_{j}})$ using Theorem 3.4. To this end let $\tau:=\mu^{(m-1)/m}\left|\det\bm{B}\right|^{(m-1)/m}.$ Inequality (4) and Theorem 3.4 give
$\operatorname{app}_{k}(\bm{A})\leq\tau\cdot\sum_{j=0}^{l-1}\left(\frac{1}{2^{\left|I_{j}\right|}}\right)^{\frac{1}{m}}\leq\tau\cdot\sum_{j=0}^{n-k-1}\left(\frac{1}{2^{n-m-j}}\right)^{\frac{1}{m}}$
since $l\leq n-k$. With the geometric sum formula one obtains
$\sum_{j=0}^{n-k-1}\left(\frac{1}{2^{n-m-j}}\right)^{\frac{1}{m}}=\frac{1}{2^{\frac{n-m}{m}}}\cdot\sum_{j=0}^{n-k-1}2^{\frac{j}{m}}=\frac{1}{2^{\frac{1}{m}}-1}\cdot\left(\frac{1}{2^{\frac{k-m}{m}}}-\frac{1}{2^{\frac{n-m}{m}}}\right).$
Proof (of Corollary 1)
The polytope $1/\left|\det\bm{B}\right|\cdot P(\bm{B})$ does not contain non-zero integer vectors in its interior since each non-zero integer vector can be expressed in terms of $\bm{B}$ and non-zero coefficients which are at least $1/\left|\det\bm{B}\right|$. Hence, if $\mu^{(m-1)/m}\cdot\left|\det\bm{B}\right|^{(2m-1)/m}<2^{(n-m)/m}$, Theorem 3.4 implies $\operatorname{app}_{\bm{B},n-1}(\bm{A})<1/\left|\det\bm{B}\right|$. This means that we can find an approximation of some $\bm{b}\in\bm{A}\mathbb{Z}^{n}_{\geq 0}$ that is contained in $\bm{b}+1/\left|\det\bm{B}\right|\cdot\operatorname{int}(P(\bm{B}))$. Since $1/\left|\det\bm{B}\right|\cdot\operatorname{int}(P(\bm{B}))$ does not contain non-zero integer vectors, the approximation error has to be zero. ∎
5 Proof of Theorem 3.3
Let $n\geq 3$. We abbreviate the bound in Theorem 3.3 by
$\epsilon:=\frac{\varphi(n-2)}{2\varphi(n-2)+1}\cdot\bm{a}_{1}.$
We first argue that without loss of generality we can suppose that $b:=\bm{a}_{2}+\ldots+\bm{a}_{n}$. If $b=\lambda_{1}\bm{a}_{1}+\ldots+\lambda_{n}\bm{a}_{n}$ for $\lambda_{i}\in\mathbb{N}$, we can assume that $\lambda_{i}>0$ for all $i\geq 2$ since, otherwise, we pass to a setting with fewer integers. Similarly, if $\lambda_{i}>1$ for an index $i\geq 2$, replace the integer $\bm{a}_{i}$ by $\tilde{\bm{a}}_{i}:=\lambda_{i}\bm{a}_{i}$. Since we fix $\bm{a}_{1}$ in our non-negative combination, we also assume that $\lambda_{1}=0$.
Let $\bm{a}_{i}=z_{i}\bm{a}_{1}+r_{i}$ with $r_{i}\in(-\bm{a}_{1}/2,\bm{a}_{1}/2]$ and $z_{i}\in\mathbb{N}$ for $i\in\{2,\ldots,n\}$. Our target is to approximate $\bm{a}_{2}+\ldots+\bm{a}_{n}$ using $\bm{a}_{1}$ and at most one additional integer $\bm{a}_{i}$. To do so, we work with the residues $r_{i}$ and approximate the sum of residues $r_{2}+\ldots+r_{n}$ using only one $r_{i}$. Without loss of generality we can assume that $r_{2}+\ldots+r_{n}>0$. Define $I:=\{i\in\{2,\ldots,n\}:r_{i}\geq 0\}$ and, hence, $r_{2}+\ldots+r_{n}\leq\sum_{i\in I}r_{i}$. We define $\tilde{s}_{i}:=\lfloor\sum_{j\in I}\bm{a}_{j}/\bm{a}_{i}\rfloor$ for all $i\in I$. Consider the following integer knapsack problem:
$\max\sum_{i\in I}r_{i}x_{i}\ \text{ s.t. }\ \sum_{i\in I}\frac{1}{\tilde{s}_{i}+1}x_{i}<1,\ x_{i}\in\mathbb{Z}_{\geq 0}\text{ for all }i\in I.$ (5)
The knapsack constraint is chosen such that the all-ones vector is a feasible solution to the problem. This is true since
$\sum_{i\in I}\frac{1}{\tilde{s}_{i}+1}<\sum_{i\in I}\frac{\bm{a}_{i}}{\sum_{j\in I}\bm{a}_{j}}=1.$
Let $\bm{z}^{*}$ be an optimal solution for (5) and $\tilde{s}_{i^{*}}r_{i^{*}}=\max_{i\in I}\tilde{s}_{i}r_{i}$. From a result of Kohli and Krishnamurti [14, Theorem 2], one obtains
$\sum_{i\in I}r_{i}\bm{z}^{*}_{i}\leq\left(1+\varphi(|I|-1)\right)\tilde{s}_{i^{*}}r_{i^{*}}.$
Since the all-ones vectors is a feasible integral solution to (5), we conclude that
$\sum_{i\in I}r_{i}-\tilde{s}_{i^{*}}r_{i^{*}}\leq\sum_{i\in I}r_{i}\bm{z}^{*}_{i}-\tilde{s}_{i^{*}}r_{i^{*}}\leq\varphi(|I|-1)\tilde{s}_{i^{*}}r_{i^{*}}.$ (6)
We next derive an upper bound on $\tilde{s}_{i^{*}}r_{i^{*}}$.
Let $s_{i^{*}}:=\lfloor\sum_{j=2}^{n}\bm{a}_{j}/\bm{a}_{i^{*}}\rfloor$ and denote by $\bar{b}$ the least positive value in $r_{2}+\ldots+r_{n}+\bm{a}_{1}\mathbb{Z}=b+\bm{a}_{1}\mathbb{Z}$. Suppose that there exists an integer $k\in\{0,\ldots,s_{i^{*}}\}$ such that $kr_{i^{*}}\in[\bar{b}-\epsilon,\bar{b}+\epsilon]$. This implies that there exists an integer $\lambda$ such that $|\lambda\bm{a}_{1}+k\bm{a}_{i^{*}}-b|=|\bar{b}-kr_{i^{*}}|\leq\epsilon$. The integer $\lambda$ is non-negative because $k\bm{a}_{i^{*}}\leq\bm{a}_{2}+\ldots+\bm{a}_{n}=b$ by definition of $s_{i^{*}}$. Hence, the theorem is correct under this assumption.
From now on we assume that $kr_{i^{*}}\notin[\bar{b}-\epsilon,\bar{b}+\epsilon]$ for all $k\in\{0,\ldots,s_{i^{*}}\}$. We claim that we can strengthen this assumption to $kr_{i^{*}}<\bar{b}-\epsilon$ for all $k\in\{0,\ldots,s_{i^{*}}\}$. Suppose this is not true. Let $k^{*}\in[s_{i^{*}}]$ be the smallest index such that $k^{*}r_{i^{*}}>\bar{b}+\epsilon$. We obtain
$r_{i^{*}}=k^{*}r_{i^{*}}-(k^{*}-1)r_{i^{*}}>\bar{b}+\epsilon-(\bar{b}-\epsilon)=2\epsilon\geq\frac{1}{2}\bm{a}_{1},$
where we use $\epsilon\geq 1/4\bm{a}_{1}$ and $(k^{*}-1)r_{i^{*}}<\bar{b}-\epsilon$ due to the minimality of $k^{*}$ and $(k^{*}-1)r_{i^{*}}\notin[\bar{b}-\epsilon,\bar{b}+\epsilon]$. This contradicts that $r_{i^{*}}\leq 1/2\bm{a}_{1}$. Thus, we can assume that $kr_{i^{*}}<\bar{b}-\epsilon$ for all $k\in\{0,\ldots,s_{i^{*}}\}$.
Since $\tilde{s}_{i^{*}}=\lfloor\sum_{j\in I}\bm{a}_{j}/\bm{a}_{i}^{*}\rfloor\leq\lfloor\sum_{j=2}^{n}\bm{a}_{j}/\bm{a}_{i}^{*}\rfloor=s_{i^{*}}$ we obtain the relation
$\tilde{s}_{i^{*}}r_{i^{*}}<\bar{b}-\epsilon.$ (7)
By definition, we have $\bar{b}<\bm{a}_{1}$. We claim that this bound can be strengthened to $\bar{b}<\bm{a}_{1}-\epsilon$ since, if $\bar{b}\geq\bm{a}_{1}-\epsilon$, we can approximate $b$ with an error of size at most $\epsilon$ by using $\lceil b/\bm{a}_{1}\rceil\bm{a}_{1}$. Inserting (7) and $\bar{b}<\bm{a}_{1}-\epsilon$ into (6) results in
$\sum_{i\in I}r_{i}-\tilde{s}_{i^{*}}r_{i^{*}}\leq\varphi(|I|-1)\tilde{s}_{i^{*}}r_{i^{*}}<\varphi(|I|-1)\left(\bar{b}-\epsilon\right)<\varphi(|I|-1)(\bm{a}_{1}-2\epsilon).$
Note that $\varphi(|I|-1)\leq\varphi(n-2)$ as $|I|\leq n-1$. So we obtain
$\sum_{i\in I}r_{i}-\tilde{s}_{i^{*}}r_{i^{*}}<\varphi(n-2)(\bm{a}_{1}-2\epsilon)=\varphi(n-2)\left(1-\frac{2\varphi(n-2)}{2\varphi(n-2)+1}\right)\bm{a}_{1}=\epsilon.$
Since $\bar{b}\leq\sum_{i\in I}r_{i}$, we deduce
$\bar{b}-\tilde{s}_{i^{*}}r_{i^{*}}\leq\sum_{i\in I}r_{i}-\tilde{s}_{i^{*}}r_{i^{*}}<\epsilon.$
However, from (7), we get $\tilde{s}_{i^{*}}r_{i^{*}}<\bar{b}-\epsilon$, which implies $\epsilon<\bar{b}-\tilde{s}_{i^{*}}r_{i^{*}}$. This contradicts the inequality above and proves the upper bound on $\operatorname{app}_{\bm{a}_{1},2}(\bm{a})$.∎
6 Lower bound constructions
We rely on the following number theoretic property.
Lemma 3
Let $\lambda_{0},\ldots,\lambda_{n-1}\in\mathbb{N}$ be given such that $\sum_{i=0}^{n-1}\lambda_{i}2^{i}\equiv 2^{n}-1\mod 2^{n}$ and $\sum_{i=0}^{n-1}\lambda_{i}\leq n$. Then we have $\lambda_{i}=1$ for all $i\in\{0,\ldots,n-1\}$.
Proof
We have that $\sum_{i=0}^{n-2}\lambda_{i}2^{i}+\lambda_{n-1}2^{n-1}$ is either $\sum_{i=0}^{n-2}\lambda_{i}2^{i}$ or $\sum_{i=0}^{n-2}\lambda_{i}2^{i}+2^{n-1}$ modulo $2^{n}$ as $2\cdot 2^{n-1}\equiv 0\mod 2^{n}$. Thus, we are only concerned with whether $\lambda_{n-1}$ is odd or even. As for the remaining coefficients, let us initially assume that $\lambda_{i}\in\{0,1\}$ for all $i\in\{0,\ldots n-2\}$. This new assumption implies $\sum_{i=0}^{n-2}\lambda_{i}2^{i}\leq 2^{n-1}-1$ by the geometric sum formula, where equality holds if and only if $\lambda_{i}=1$ for all $i\in\{0,\ldots,n-2\}$. Since $\sum_{i=0}^{n-1}\lambda_{i}2^{i}\equiv 2^{n}-1\mod 2^{n}$, it follows that we are in the case where $\lambda_{n-1}$ is odd, that is, $\sum_{i=0}^{n-2}\lambda_{i}2^{i}+2^{n-1}$. So we have $\lambda_{i}=1$ for all $i\in\{0,\ldots,n-2\}$ and $\lambda_{n-1}$ being odd. As $\lambda_{n-1}\leq n-\sum_{i=0}^{n-2}\lambda_{i}=1$, we conclude that $\lambda_{n-1}=1$ as well.
Let at least one of the $\lambda_{i}$ for $i\in\{0,\ldots,n-2\}$ not be in $\{0,1\}$. We will show that this implies $n<\sum_{i=0}^{n-1}\lambda_{i}$, which proves the lemma since the previous case is the unique solution to $\sum_{i=0}^{n-1}\lambda_{i}2^{i}\equiv 2^{n}-1\mod 2^{n}$ with $\sum_{i=0}^{n-1}\lambda_{i}\leq n$. Let $\lambda_{i}\geq 2$ for some $i\in\{0,\ldots,n-2\}$. We reallocate the coefficients using
$\lambda_{i}2^{i}=2^{i+1}+(\lambda_{i}-2)2^{i}.$
This operation shrinks the coefficient of $2^{i}$ by two while simultaneously decreasing the coefficient sum by one. We apply this operation until the new coefficients satisfy $\tilde{\lambda}_{i}\in\{0,1\}$ for all $i\in\{0,\ldots,n-2\}$. So we are in the setting from above. There, the unique solution is given when $\tilde{\lambda}_{i}=1$ for all $i\in\{0,\ldots,n-1\}$ and $\sum_{i=0}^{n-1}\tilde{\lambda}_{i}=n$. However, since our operation strictly decreases the coefficient sum in each step, we have $n=\sum_{i=0}^{n-1}\tilde{\lambda}_{i}<\sum_{i=0}^{n-1}\lambda_{i}$. ∎
We begin by proving that the bounds in Theorem 3.1 are tight.
Proposition 1
Let $n\geq 2$. There exists $\bm{a}\in\mathbb{Z}^{n}$ with $0<\bm{a}_{1}\leq\bm{a}_{2}\leq\ldots\leq\bm{a}_{n}$ such that
$\frac{1}{2^{n-1}}\cdot\bm{a}_{1}\leq\operatorname{app}_{n-1}(\bm{a}).$
Proof
Choose $\bm{a}_{1}:=2^{n-1}$ and $\bm{a}_{i}:=2^{n}+2^{i-2}$ for $i\in\{2,\ldots,n\}$. We want to approximate $b:=\sum_{i=1}^{n}\bm{a}_{i}=2^{n}n-1$. If $\bar{b}\in\mathbb{N}$ is a possible approximation with $b\neq\bar{b}$, we have $\left|b-\bar{b}\right|\geq 1=1/2^{n-1}\cdot\bm{a}_{1}$. So it suffices to verify that it is impossible to generate $b$ using $n-1$ integers $\bm{a}_{1},\ldots,\bm{a}_{n}$. Let $\lambda_{1},\ldots,\lambda_{n}\in\mathbb{N}$ be such that $\sum_{i=1}^{n}\lambda_{i}\bm{a}_{i}=b$. We have $\sum_{i=2}^{n}\lambda_{i}\leq n-1$, because otherwise we get $\sum_{i=2}^{n}\lambda_{i}\bm{a}_{i}>2^{n}n>\bm{b}$. Since $\sum_{i=1}^{n}\lambda_{i}\bm{a}_{i}=\bm{b}$, we also have $\sum_{i=1}^{n}\lambda_{i}\bm{a}_{i}\equiv\bm{b}\mod 2^{n}$. This is equivalent to $\tilde{\lambda}_{1}2^{n-1}+\sum_{i=2}^{n}\lambda_{i}2^{i-2}\equiv 2^{n}-1\mod 2^{n}$, where $\tilde{\lambda}_{1}\in\{0,1\}$ and $\tilde{\lambda}_{1}+\sum_{i=2}^{n}\lambda_{i}\leq n$. By Lemma 3, we conclude that $\lambda_{i}=1$ for all $i\in\{2,\ldots,n\}$ and $\tilde{\lambda}_{1}=1$. Thus, generating $\bm{b}$ using $\bm{a}_{1},\ldots,\bm{a}_{n}$ requires $\lambda_{i}\geq 1$ for all $i\in[n]$. ∎
Next, we show that our bounds in Theorem 3.3 are tight as well.
Proposition 2
Let $n\geq 2$. There exists $\bm{a}\in\mathbb{Z}^{n}$ with $0<\bm{a}_{1}\leq\bm{a}_{2}\leq\ldots\leq\bm{a}_{n}$ such that
$\frac{\varphi(n-2)}{2\varphi(n-2)+1}\cdot\bm{a}_{1}\leq\operatorname{app}_{2}(\bm{a}).$
Proof
For $n=2$, the claim follows directly since $\varphi(n-2)$ is the empty sum which equals 00. When $n=3$, we apply the tight bound from Proposition 1 and observe that $\varphi(1)=1/2$, which implies $\varphi(1)/(2\varphi(1)+1)=1/4$. So let $n\geq 4$. Observe that $\operatorname{app}_{2}(\bm{a})=\operatorname{app}_{2}(\kappa\cdot\bm{a})$ for some scalar $\kappa>0$. The same holds for $\operatorname{app}_{\bm{a}_{1},2}(\bm{a})$. So we assume without loss of generality that $\bm{a}_{1}:=1$ and construct a rational vector $\bm{a}\in\mathbb{Q}^{n}$. The integrality follows then by scaling with the least common multiple of the denominators. Recall that $\{t_{i}+1\}_{i\in\mathbb{N}}$ denotes the Sylvester sequence. Set $r_{n}:=1/(2\varphi(n-2)+1)$ and define
$z_{i}:=\prod_{\begin{subarray}{c}j=0,\\ j\neq n-i\end{subarray}}^{n-2}(t_{j}+1)\text{ and }r_{i}:=\frac{1}{t_{n-i}}\cdot r_{n}$
for $i\in\{2,\ldots,n\}$. Let $\tau>0$ be a fixed sufficiently large natural number. It will become clear from the arguments below what sufficient means in this context. We define $\bm{a}_{i}:=\tau\cdot z_{i}+r_{i}$ and try to approximate $b:=\bm{a}_{2}+\ldots+\bm{a}_{n}$. First, we approximate $b$ using $\bm{a}_{1}$ and one more $\bm{a}_{i}$ similar to the proof of Theorem 3.3. We define $s_{i}:=\lfloor\sum_{j=2}^{n}\bm{a}_{j}/\bm{a}_{i}\rfloor$. So our candidates to approximate $\bm{a}_{2}+\ldots+\bm{a}_{n}$ from below are $\mu\bm{a}_{i}+\lambda\bm{a}_{1}$ with some suitable $\lambda\in\mathbb{N}$ for $\mu\in\{0,\ldots,s_{i}\}$ and $i\in\{2,\ldots,n\}$. If $\mu=0$, we approximate $b$ using only $\bm{a}_{1}=1$. This is equivalent to rounding $r_{2}+\ldots+r_{n}$ to the nearest integer. We have
$r_{2}+\ldots+r_{n}=r_{n}\sum_{i=2}^{n}\frac{1}{t_{n-i}}=r_{n}(\varphi(n-2)+1)=\frac{\varphi(n-2)+1}{2\varphi(n-2)+1}.$
Rounding this expression to the nearest integer gives us precisely an approximation error of $\varphi(n-2)/(2\varphi(n-2)+1)$. To improve upon this bound, we have to incorporate one of the other $\bm{a}_{i}$. For that, we need an upper bound on $s_{i}$ to determine how often we can use each $\bm{a}_{i}$. We have
$s_{i}=\left\lfloor\frac{\sum_{j=2}^{n}\bm{a}_{j}}{\bm{a}_{i}}\right\rfloor=\left\lfloor\frac{\tau\sum_{j=2}^{n}z_{j}}{\tau z_{i}+r_{i}}+\frac{\sum_{j=2}^{n}r_{j}}{\tau z_{i}+r_{i}}\right\rfloor\leq\left\lfloor\frac{\sum_{j=2}^{n}z_{j}}{z_{i}}+\frac{1}{\tau}c\right\rfloor,$
where we abbreviate $c=\sum_{j=2}^{n}r_{j}/z_{i}$. We continue the calculation and utilize the identity $\sum_{j=0}^{n-2}1/(t_{j}+1)=1-1/(t_{n-1}+1)$ from [21] to obtain
$s_{i}\leq\left\lfloor\frac{\sum_{j=2}^{n}z_{j}}{z_{i}}+\frac{1}{\tau}c\right\rfloor$ $=\left\lfloor(t_{n-i}+1)\cdot\sum_{j=2}^{n}\frac{1}{t_{n-j}+1}+\frac{1}{\tau}c\right\rfloor$
$=\left\lfloor(t_{n-i}+1)\cdot\sum_{j=0}^{n-2}\frac{1}{t_{j}+1}+\frac{1}{\tau}c\right\rfloor$
$=\left\lfloor(t_{n-i}+1)\cdot\left(1-\frac{1}{t_{n-1}+1}\right)+\frac{1}{\tau}c\right\rfloor$
$=\left\lfloor t_{n-i}+1-\frac{t_{n-i}+1}{t_{n-1}+1}+\frac{1}{\tau}c\right\rfloor=t_{n-i},$
where we use $t_{n-i}<t_{n-1}$ for all $i\in\{2,\ldots,n\}$ and that $\tau$ is sufficiently large in the last step. We established $s_{i}\leq t_{n-i}$ for all $i\in\{2,\ldots,n\}$. This implies $s_{i}r_{i}\leq r_{n}$ by the definition of $r_{i}$ for all $i\in\{2,\ldots,n\}$. We get
$r_{2}+\ldots+r_{n}-s_{i}r_{i}\geq r_{2}+\ldots r_{n}-r_{n}=r_{2}+\ldots+r_{n-1}.$
So the best approximation of $b$ using $\mu\bm{a}_{i}+\lambda\bm{a}_{1}$ with $\mu\geq 1$ gives us
$\left|b-(\mu\bm{a}_{i}+\lambda\bm{a}_{1})\right|\geq\left|r_{2}+\ldots+r_{n}-s_{i}r_{i}\right|\geq\left|r_{2}+\ldots+r_{n-1}\right|.$
The approximation error equals
$r_{2}+\ldots+r_{n-1}=r_{n}\cdot\sum_{j=2}^{n-1}\frac{1}{t_{n-j}}=\frac{\varphi(n-2)}{2\varphi(n-2)+1}.$
This is again our desired lower bound. So one cannot improve upon this bound by using $\bm{a}_{1}$ and one of the other $\bm{a}_{i}$. As a last step, we ensure that there is no other integral non-negative combination of at most two integers $\bm{a}_{i}$ and $\bm{a}_{j}$ where none of them equals $\bm{a}_{1}$. Let $\lambda_{i},\lambda_{j}\in\mathbb{N}$. Then we have
$\left|b-\lambda_{i}\bm{a}_{i}-\lambda_{j}\bm{a}_{j}\right|=\left|\tau\left(\sum_{j=2}^{n}z_{j}-\lambda_{i}z_{i}-\lambda_{j}z_{j}\right)+\sum_{j=2}^{n}r_{j}-\lambda_{i}r_{i}-\lambda_{j}r_{j}\right|.$
We abbreviate the last term with $\left|\tau d+e\right|$. We have $d\neq 0$ since $\lambda_{i}z_{i}+\lambda_{j}z_{j}$ is divisible by $(t_{k}+1)$ for $t_{k}\notin\{t_{n-i},t_{n-j}\}$ but $\sum_{j=2}^{n}z_{j}$ is not. This implies that the error grows as $\tau$ becomes larger. This holds for all possible coefficients $\lambda_{i}$ and $\lambda_{j}$. Recall that we have $\lambda_{i}\leq s_{i}\leq t_{n-i}$ for all $i\in\{2,\ldots,n\}$ if $\tau$ is large enough. So we only need to consider finitely many of $(\lambda_{i},\lambda_{j})$ combinations for $i,j\in\{2,\ldots,n\}$. Hence, by choosing $\tau$ large enough, we increase the approximation error for any combination that can be formed without using $\bm{a}_{1}$. This shows that $\varphi(n-2)/(2\varphi(n-2)+1)\cdot\bm{a}_{1}\leq\operatorname{app}_{2}(\bm{a})$. ∎
We show that for $m\geq 2$ it is necessary to incorporate additional information from the matrix to obtain a decreasing behavior in $k$.
Proposition 3
Let $n\geq 3$. There exists $\bm{A}=(\bm{a}_{1},\ldots,\bm{a}_{n})\in\mathbb{Z}^{2\times n}$ such that $\bm{a}_{i}\in\operatorname{pos}\{\bm{a}_{1},\bm{a}_{2}\}$ for all $i\in[n]$ and
$\frac{\left\lfloor\sqrt{n-1}\right\rfloor}{n-1}\leq\operatorname{app}_{k}(\bm{A})$
for all $k\in\{2,\ldots,n-1\}$.
Proof
Let $q:=\lfloor\sqrt{n-1}\rfloor$. Define $\bm{a}_{1}:=(n-1)\bm{e}_{1}$, $\bm{a}_{2}:=(n-1)\bm{e}_{2}$, $\bm{B}:=(\bm{a}_{1},\bm{a}_{2})$, and $\bm{r}:=(1,q)^{\top}$, where $\bm{e}_{1},\bm{e}_{2}$ are the two canonical unit vectors. Set $\bm{a}_{i}:=(n-1)\bm{z}_{i}+\bm{r}$ with
$\bm{z}_{i}:=(2^{n-2}+1)\begin{pmatrix}1\\ 1\end{pmatrix}+2^{i-2}\begin{pmatrix}1\\ -1\end{pmatrix}$
for all $i\in\{3,\ldots,n\}$. The coefficient $12^{n-2}+$ 1 is chosen such that each component of $\bm{z}_{i}$ is positive for all $i\in\{3,\ldots,n\}$. This guarantees $\bm{a}_{i}\in\operatorname{pos}\{\bm{a}_{1},\bm{a}_{2}\}$ for all $i\in\{3,\ldots,n\}$. We aim to approximate $\bm{b}:=\bm{a}_{1}+\ldots+\bm{a}_{n}$ with some $\tilde{\bm{b}}:=\lambda_{1}\bm{a}_{1}+\ldots+\lambda_{n}\bm{a}_{n}$ for $\lambda_{1},\ldots,\lambda_{n}\in\mathbb{N}$, where $n-k$ of the $\lambda_{i}$ are zero. We distinguish between the cases where $\sum_{i=3}^{n}\lambda_{i}\leq n-3$, $\sum_{i=3}^{n}\lambda_{i}=n-2$, and $\sum_{i=3}^{n}\lambda_{i}\geq n-1$. It will turn out that in the case $\sum_{i=3}^{n}\lambda_{i}\leq n-3$ the approximation can never improve upon $q/(n-1)$. In the other two cases, we will show that the approximation $\tilde{\bm{b}}$ is by construction already too far away from $\bm{b}$.
Suppose that $\sum_{i=3}^{n}\lambda_{i}\leq n-3$. We check the approximation error modulo $\bm{B}\mathbb{Z}^{2}$. For $\bm{y}\in\mathbb{R}^{2}$, let $\bar{\bm{y}}$ denote the unique vector from $\bm{y}+(\bm{a}_{1},\bm{a}_{2})\mathbb{Z}^{2}$ that is contained in the half-open parallelepiped $(n-1)[0,1)^{2}$. In particular, we have $\overline{\bm{b}}=\overline{(n-2)\bm{r}}$ and $\overline{\tilde{\bm{b}}}=\overline{(\lambda_{3}+\ldots+\lambda_{n})\bm{r}}$. We claim that the choice of $\bm{r}$ implies
$\|\overline{(n-2)\bm{r}}-\overline{\lambda\bm{r}}\|_{P(\bm{B})}\geq\frac{q}{n-1}$ (8)
for all $\lambda\in\{0,\ldots,n-3\}$. In fact, we will show a slightly stronger statement. Let $\Lambda:=(\bm{a}_{1},\bm{a}_{2},\bm{r})\mathbb{Z}^{3}$. We prove that, if $\bm{x},\bm{y}\in\Lambda$ with $\bm{x}\notin\bm{y}+\bm{B}\mathbb{Z}^{2}$, then $\|\bm{x}-\bm{y}\|_{\infty}\geq q$. This implies the claim above since $\overline{(n-2)\bm{r}}\notin\overline{\lambda\bm{r}}+\bm{B}\mathbb{Z}^{2}$ and $\|\bm{x}\|_{P(\bm{B})}=1/(n-1)\|\bm{x}\|_{\infty}$ for all $\bm{x}\in\mathbb{R}^{2}$. First, observe that minimizing $\|\bm{x}-\bm{y}\|_{\infty}$ over $\bm{x},\bm{y}\in\Lambda$ with $\bm{x}\notin\bm{y}+\bm{B}\mathbb{Z}^{2}$ is equivalent to minimizing $\|\bm{z}\|_{\infty}$ over $\bm{z}\in\Lambda\backslash\{\bm{0}\}$. In particular, we can focus on $\bm{z}\in\Lambda\backslash\bm{B}\mathbb{Z}^{2}$. Those lattice vectors are of the form $t\bm{r}+\bm{B}\mathbb{Z}^{2}$ for $t\in\{1,\ldots,n-2\}$. To improve upon (8), both coordinates of $\bm{z}$ have to be at most $q-1$ in absolute value. Since $\bm{r}_{1}=1$, this is only possible if $t\in[q-1]\cup\{n-1-(q-1),\ldots,n-2\}$. By symmetry, we focus without loss of generality on the case when $t\in[q-1]$. Analyzing the second coordinate, $\bm{r}_{2}=q$, we get for $j\in[q-1]$ that
$q\leq j\bm{r}_{2}\leq q^{2}-q\leq n-1-q.$
This shows $\|\bm{z}\|_{\infty}\geq q$ for all $\bm{z}\in\Lambda\backslash\{\bm{0}\}$. The inequality in (8) follows. Since adding multiples of $\bm{a}_{1}$ and/or $\bm{a}_{2}$ does not affect the estimate in (8), we deduce that, under the assumption that $\sum_{i=3}^{n}\lambda_{i}\leq n-3$, we cannot achieve an approximation better than $q/(n-1)$.
Let us next assume that $\sum_{i=3}^{n}\lambda_{i}\geq n-1$. By construction, the vectors $\bm{a}_{3},\ldots,\bm{a}_{n}$ satisfy $\bm{1}^{\top}\bm{a}_{i}=\bm{1}^{\top}\bm{a}_{j}$ for $i,j\in\{3,\ldots,n\}$. We use this fact to measure the distance between $\bm{b}$ and $\tilde{\bm{b}}$ as follows:
$\bm{1}^{\top}\tilde{\bm{b}}-\bm{1}^{\top}\bm{b}\geq\bm{1}^{\top}(-\bm{a}_{1}-\bm{a}_{2})+\bm{1}^{\top}\bm{a}_{3}\sum_{i=3}^{n}(\lambda_{i}-1)\geq-2(n-1)+\bm{1}^{\top}\bm{a}_{3},$
where in the final inequality we utilize the assumption $\sum_{i=3}^{n}\lambda_{i}\geq n-1$. Next we compute
$\bm{1}^{\top}\bm{a}_{3}=(n-1)\bm{1}^{\top}\bm{z}_{3}+\bm{1}^{\top}\bm{r}=2^{n-1}(n-1)+2(n-1)+q+1,$
which gives us $\bm{1}^{\top}\tilde{\bm{b}}-\bm{1}^{\top}\bm{b}\geq 2^{n-1}(n-1)+q+1>2(n-1)$ as $n\geq 3$. We conclude with
$\|\tilde{\bm{b}}-\bm{b}\|_{P(\bm{B})}=\frac{1}{n-1}\|\tilde{\bm{b}}-\bm{b}\|_{\infty}\geq\frac{1}{2(n-1)}\bm{1}^{\top}(\tilde{\bm{b}}-\bm{b})>1$
using that $2\|\bm{x}\|_{\infty}\geq\|\bm{x}\|_{1}\geq\bm{1}^{\top}\bm{x}$ for all $\bm{x}\in\mathbb{R}^{2}$. This shows that $\tilde{\bm{b}}$ is too far away to be a decent approximation.
We are left with the case $\sum_{i=3}^{n}\lambda_{i}=n-2$. The vector $\tilde{\bm{b}}-(\lambda_{1}\bm{a}_{1}+\lambda_{2}\bm{a}_{2})$ equals
$(n-1)\left((2^{n-2}+1)(n-2)\begin{pmatrix}1\\ 1\end{pmatrix}+\left(\sum_{i=3}^{n}\lambda_{i}2^{i-2}\right)\begin{pmatrix}1\\ -1\end{pmatrix}\right)+(n-2)\bm{r}$
and $\bm{b}$ equals
$\bm{a}_{1}+\bm{a}_{2}+(n-1)\left((2^{n-2}+1)(n-2)\begin{pmatrix}1\\ 1\end{pmatrix}+2\left(2^{n-2}-1\right)\begin{pmatrix}1\\ -1\end{pmatrix}\right)+(n-2)\bm{r}$
since the identity $\sum_{i=3}^{n}2^{i-2}=2(2^{n-2}-1)$ holds. We get the following difference
$\tilde{\bm{b}}-\bm{b}=(\lambda_{1}-1)\bm{a}_{1}+(\lambda_{2}-1)\bm{a}_{2}+2(n-1)\underbrace{\left(\sum_{i=3}^{n}\lambda_{i}2^{i-3}-(2^{n-2}-1)\right)}_{=:d}\begin{pmatrix}1\\ -1\end{pmatrix}.$
Suppose that $d\neq 0$. We have $d\in\mathbb{Z}$ and assume that $d\geq 1$. Then the first coordinate of $\tilde{\bm{b}}-\bm{b}$ satisfies
$(\lambda_{1}-1)(n-1)+2(n-1)d\geq n-1.$
Therefore, the approximation error can be estimated by $\|\bm{b}-\tilde{\bm{b}}\|_{P(\bm{B})}\geq 1$. If $d\leq-1$, the same argument applies to the second coordinate. We are left with the case when $d=0$. Under this assumption, we have to have $\lambda_{1}=1=\lambda_{2}$, otherwise one coordinate is again at least $n-1$ in absolute value. However, if $\lambda_{1}=1=\lambda_{2}$ and $d=0$, we get $\lambda_{i}=1$ for all $i\in[n]$ by Lemma 3. ∎